# ButterflyQuant: Ultra-low-bit LLM Quantization through Learnable Orthogonal Butterfly Transforms

### v1

@@ Abstract
Large language models require massive memory footprints, severely limiting deployment on consumer hardware. Quantization reduces memory through lower numerical precision, but extreme 2-bit quantization suffers from catastrophic performance loss due to outliers in activations. Rotation-based methods such as QuIP and QuaRot apply orthogonal transforms to eliminate outliers before quantization, using computational invariance: $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$ for orthogonal $\mathbf{Q}$. However, these methods use fixed transforms–Hadamard matrices achieving optimal worst-case coherence $\mu=1/\sqrt{n}$–that cannot adapt to specific weight distributions. We identify that different transformer layers exhibit distinct outlier patterns, motivating layer-adaptive rotations rather than one-size-fits-all approaches. We propose ButterflyQuant, which replaces Hadamard rotations with learnable butterfly transforms parameterized by continuous Givens rotation angles. Unlike Hadamard’s discrete $\{+1,-1\}$ entries that are non-differentiable and prohibit gradient-based learning, butterfly transforms’ continuous parameterization enables smooth optimization while guaranteeing orthogonality by construction. This orthogonal constraint ensures theoretical guarantees in outlier suppression while achieving $O(n\log n)$ computational complexity with only $\frac{n\log n}{2}$ learnable parameters. We further introduce a uniformity regularization on post-transformation activations to promote smoother distributions amenable to quantization. Learning requires only 128 calibration samples and converges in minutes on a single GPU–a negligible one-time cost. On LLaMA-2-7B with 2-bit quantization, ButterflyQuant achieves 15.4 perplexity versus 22.1 for QuaRot.
@@ 1 Introduction
Figure 1: Layer heterogeneity motivates learnable transforms for LLM quantization. (a) Different transformer layers exhibit distinct outlier distributions: attention (positive tails), early MLP (negative regions), late MLP (boundaries). (b) Hadamard transforms with discrete $\{+1,-1\}$ entries apply fixed rotations through recursive decomposition $\mathbf{H}_{2n}=\frac{1}{\sqrt{2}}[\mathbf{H}_{n},\mathbf{H}_{n};\mathbf{H}_{n},-\mathbf{H}_{n}]$, achieving uniform coherence $\mu=1/\sqrt{n}=0.0156$ across all layers. (c) Butterfly transforms use continuous rotation angles $\theta_{i,j}$ in Givens rotations $\mathbf{G}(\theta)$, enabling gradient-based optimization to learn layer-specific patterns. This yields adaptive coherence (e.g., $\mu\in[0.018,0.032]$) that matches each layer’s unique outlier distribution.
@@
To address this need for an adaptive yet efficient orthogonal transform, we propose ButterflyQuant. Our method replaces fixed Hadamard rotations with learnable butterfly transforms–structured orthogonal matrices factorized into $O(n\log n)$ Givens rotations [Dao et al., 2019]. As shown in Figure 1(c), butterfly transforms maintain the efficient computational structure of Hadamard matrices but learn layer-specific rotation angles via gradient descent. Crucially, butterfly transforms parameterize rotations through continuous angles $\theta\in\mathbb{R}$, enabling smooth gradient flow and stable optimization—in contrast to Hadamard’s discrete $\{+1,-1\}$ entries that prohibit gradient-based learning. This allows them to develop adaptive coherence patterns that are tailored to each layer’s unique outlier distribution. Unlike other learnable methods (e.g., SpinQuant [Liu et al., 2024a]) that optimize over the full Stiefel manifold with high computational cost, our sparse parameterization guarantees orthogonality by construction, enabling stable and efficient optimization. For non-power-of-2 dimensions common in LLMs (e.g., 5120), we develop composite transforms using Kronecker products, extending our method’s applicability. The learning process is remarkably lightweight, converging in minutes on a single GPU with a small calibration set, making it a practical and effective solution.
@@ 2.1 Post-Training Quantization for LLMs
Compensation and Salience Methods. GPTQ [Frantar et al., 2023] uses second-order Hessian information for reconstruction error minimization. AWQ [Lin et al., 2024] preserves salient weights while quantizing others based on activation magnitudes. SmoothQuant [Xiao et al., 2023] migrates difficulty from activations to weights via channel-wise scaling. OmniQuant [Shao et al., 2023] and AffineQuant [Guan et al., 2024a] apply learnable transformations for outlier suppression. Mixed-precision methods like LLM.int8() [Dettmers et al., 2022] and SpQR [Dettmers et al., 2023] maintain outlier channels in higher precision but increase complexity.
@@ 2.2 Rotation-Based Quantization
Learned Methods. SpinQuant [Liu et al., 2024a] optimizes full $n\times n$ rotation matrices on the Stiefel manifold, requiring $O(n^{2})$ parameters (16.7M for 4096-dim layers) and $O(n^{3})$ operations. ROSAQ [Kim et al., 2024] and KurTail [Li et al., 2025] learn saliency-aware and kurtosis-guided rotations respectively but lack theoretical guarantees and still require $O(n^{2})$ space.
@@ 2.3 Structured Transforms and Orthogonal Parameterization
Despite extensive use in deep learning, structured transforms remain unexplored for quantization, which requires: (1) strict orthogonality for computational invariance $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$, (2) adaptability to layer-specific outliers [Sun et al., 2024], and (3) efficient inference. While FlatQuant [Liu et al., 2024c] uses Kronecker decomposition, it sacrifices orthogonality. ButterflyQuant uniquely combines continuous parameterization ($\theta\in\mathbb{R}$ enabling gradients vs. Hadamard’s discrete $\{+1,-1\}$) and $O(n\log n)$ complexity balancing expressiveness with efficiency -the first to leverage these properties for LLM quantization.
@@ 3.3.2 Concrete Example for d = 5120
where $\mathbf{A}$ is skew-symmetric ($\mathbf{A}^{T}=-\mathbf{A}$) with zeros on the diagonal, requiring $\frac{40\times(40-1)}{2}=780$ parameters. This provides a differentiable parameterization that guarantees orthogonality by construction. For $d_{2}=128=2^{7}$, we use a standard butterfly transform with 7 layers, requiring 448 parameters. The total of 1,228 parameters achieves a 21,347× reduction versus a full 5120 × 5120 Hadamard matrix.
@@ 4 Experiments
Table 1: Comprehensive evaluation of 2-bit weight quantization (W2A16) on LLaMA-2 models.
@@ 4.1 Experimental Setup
We evaluate ButterflyQuant on LLaMA-2-7B and LLaMA-2-13B [Touvron et al., 2023], using WikiText-2 [Merity et al., 2017] and C4 [Raffel et al., 2020] for perplexity evaluation, plus six zero-shot reasoning tasks: WinoGrande [Sakaguchi et al., 2020], PIQA [Bisk et al., 2020], HellaSwag [Zellers et al., 2019], ARC [Clark et al., 2018], and MMLU [Hendrycks et al., 2021]. We compare against GPTQ [Frantar et al., 2023], AWQ [Lin et al., 2024], OmniQuant [Shao et al., 2023], and QuIP [Chee et al., 2023] under aggressive 2-bit weight quantization (W2A16, denoting 2-bit weights with 16-bit activations).
@@ 4.2 Main Results
Table 1 demonstrates ButterflyQuant’s superiority in low-bit quantization across all metrics. We achieve 2.4× lower perplexity than the best baseline (15.40 vs 36.77 for GPTQ on WikiText-2 7B, 10.24 vs 13.75 for QuIP on 13B) while AWQ fails catastrophically with perplexity exceeding $10^{5}$. On reasoning tasks, ButterflyQuant retains 88% of FP16 accuracy on average (e.g., 62.27% vs 69.06% FP16 on WinoGrande, 68.97% vs 78.07% FP16 on PIQA) while baselines retain only 65-73% of FP16 performance, with consistent improvements across both model scales validating our layer-adaptive approach.
@@ 4.3 Ablation Studies
4.4 Training Dynamics and Design Validation
@@ A.1 Implementation Details
Optimization. Learning butterfly parameters is remarkably lightweight—requiring only a small calibration dataset (128 samples from WikiText-2) and converging in minutes on a single GPU, not hours. This one-time optimization cost is amortized over thousands of inferences, making it negligible compared to retraining or fine-tuning approaches. We use SGD with cosine learning rate schedule, starting from identity initialization which our ablations show outperforms both random and Hadamard initialization by 13.4% and 6.3% respectively. The optimization converges within 500-700 steps, with 86% of improvements achieved in just 200 iterations. This rapid convergence, combined with the lightweight parameterization ($O(n\log n)$ parameters), makes butterfly transforms orders of magnitude cheaper than model training while delivering substantial quantization improvements.
@@
A.2 Example: 4×4 Hadamard
@@
Appendix B Butterfly Transform Details
B.1 Concrete Example: Butterfly Matrix Structure for $n=8$
@@
B.2 The Butterfly Pattern Visualization
@@
B.3 Permutation Structure and Block Decomposition
@@
B.4 Alternative Factorizations for Non-Power-of-2 Dimensions
@@
  • $80\times 64$: More uniform but requires composite butterfly for both factors
@@ B.5 Mathematical Properties
Butterfly transforms with $O(\log n)$ layers can efficiently approximate structured orthogonal matrices and exactly represent important transforms including Hadamard, DFT, and DCT matrices.
@@
B.6 Method Characteristics and Efficiency

### v2

@@ Abstract
Large language models require massive memory footprints, severely limiting deployment on consumer hardware. Quantization reduces memory through lower numerical precision, but extreme 2-bit quantization suffers from catastrophic performance loss due to outliers in activations. Rotation-based methods such as QuIP and QuaRot apply orthogonal transforms to eliminate outliers before quantization, using computational invariance: $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$ for orthogonal $\mathbf{Q}$. However, these methods use fixed transforms–Hadamard matrices achieving optimal worst-case coherence $\mu=1/\sqrt{n}$–that cannot adapt to specific weight distributions. We identify that different transformer layers exhibit distinct outlier patterns, motivating layer-adaptive rotations rather than one-size-fits-all approaches. In this work, we propose ButterflyQuant, which replaces Hadamard rotations with learnable butterfly transforms parameterized by continuous Givens rotation angles. Unlike Hadamard’s discrete $\{+1,-1\}$ entries that are non-differentiable and thus prohibit gradient-based learning, butterfly transforms’ continuous parameterization enables smooth optimization while guaranteeing orthogonality by construction. This orthogonal constraint ensures theoretical guarantees in outlier suppression while achieving $O(n\log n)$ computational complexity with only $\frac{n\log n}{2}$ learnable parameters. We further introduce a uniformity regularization on post-transformation activations to promote smoother distributions amenable to quantization. Learning requires only 128 calibration samples and converges in minutes on a single GPU–a negligible one-time cost. For LLaMA-2-7B with 2-bit quantization, ButterflyQuant achieves 15.4 perplexity versus 37.3 for QuIP. Codes are available.
@@ 1 Introduction
Figure 1: Layer heterogeneity motivates learnable transforms for LLM quantization. (a) Different transformer layers exhibit distinct outlier distributions: attention (positive tails), early MLP (negative regions), late MLP (boundaries). (b) Hadamard transforms with discrete $\{+1,-1\}$ entries apply fixed rotations through recursive decomposition $\mathbf{H}_{2n}=\frac{1}{\sqrt{2}}[\mathbf{H}_{n},\mathbf{H}_{n};\mathbf{H}_{n},-\mathbf{H}_{n}]$, achieving uniform coherence $\mu=1/\sqrt{n}=0.0156$ across all layers. (c) Butterfly transforms use continuous rotation angles $\theta_{i,j}$ in Givens rotations $\mathbf{G}(\theta)$, enabling gradient-based optimization to learn layer-specific patterns. This yields adaptive coherence that matches each layer’s outlier distribution.
@@
To address this need for an adaptive yet efficient orthogonal transform, we propose ButterflyQuant. Our method replaces fixed Hadamard rotations with learnable butterfly transforms–structured orthogonal matrices factorized into $O(n\log n)$ Givens rotations [Dao et al., 2019]. As shown in Figure 1(c), butterfly transforms maintain the efficient computational structure of Hadamard matrices but learn layer-specific rotation angles via gradient descent. Crucially, butterfly transforms parameterize rotations through continuous angles $\theta\in\mathbb{R}$, enabling smooth gradient flow and stable optimization—in contrast to Hadamard’s discrete $\{+1,-1\}$ entries that prohibit gradient-based learning. This allows them to develop adaptive coherence patterns that are tailored to each layer’s unique outlier distribution. Unlike other learnable methods (e.g., SpinQuant [Liu et al., 2024a]) that optimize over the full Stiefel manifold with high computational cost, our sparse parameterization guarantees orthogonality by construction, enabling stable and efficient optimization. For non-power-of-2 dimensions common in LLMs (e.g., 5120), we develop composite transforms using Kronecker products. The learning process is remarkably lightweight, converging in minutes on a single GPU with a small calibration set.
@@ 2.1 Post-Training Quantization for LLMs
Compensation and Salience Methods. GPTQ [Frantar et al., 2023] uses second-order Hessian information for reconstruction error minimization. AWQ [Lin et al., 2024a] preserves salient weights while quantizing others based on activation magnitudes. SmoothQuant [Xiao et al., 2023] migrates difficulty from activations to weights via channel-wise scaling. OmniQuant [Shao et al., 2023] and AffineQuant [Guan et al., 2024a] apply learnable transformations for outlier suppression. Mixed-precision methods like LLM.int8() [Dettmers et al., 2022] and SpQR [Dettmers et al., 2023] maintain outlier channels in higher precision but increase complexity.
@@ 2.2 Rotation-Based Quantization
Learned Methods. SpinQuant [Liu et al., 2024a] optimizes full $n\times n$ rotation matrices on the Stiefel manifold, requiring $O(n^{2})$ parameters (16.7M for 4096-dim layers) and $O(n^{3})$ operations. ROSAQ [Kim et al., 2024] and KurTail [Akhondzadeh et al., 2025] learn saliency-aware and kurtosis-guided rotations respectively, but lack theoretical guarantees and still require $O(n^{2})$ space. DuQuant Lin et al. [2024b] employs a greedy construction method that uses identified outlier dimensions as prior knowledge to build rotation matrices, followed by zigzag permutation to balance outlier distribution across blocks.
@@ 2.3 Structured Transforms and Orthogonal Parameterization
Despite extensive use in deep learning, structured transforms remain unexplored for quantization, which requires: (1) strict orthogonality for computational invariance $\mathbf{y}=\mathbf{Wx}=(\mathbf{WQ}^{T})(\mathbf{Qx})$, (2) adaptability to layer-specific outliers [Sun et al., 2024], and (3) efficient inference. While FlatQuant [Liu et al., 2024c] uses Kronecker decomposition, it sacrifices orthogonality. ButterflyQuant uniquely combines continuous parameterization ($\theta\in\mathbb{R}$ enabling gradients vs. Hadamard’s discrete $\{+1,-1\}$) and $O(n\log n)$ complexity, balancing expressiveness with efficiency. ButterflyQuant is the pioneer in leveraging these properties for LLM quantization.
@@ 3.3.2 Concrete Example for d = 5120
where $\mathbf{A}$ is skew-symmetric ($\mathbf{A}^{T}=-\mathbf{A}$) with zeros on the diagonal, requiring $\frac{40\times(40-1)}{2}=780$ parameters. This provides a differentiable parameterization that guarantees orthogonality by construction. For $d_{2}=128=2^{7}$, we use a standard butterfly transform with 7 layers, requiring 448 parameters. The total of 1,228 parameters achieves a 21,347× reduction versus a full 5120 × 5120 matrix.
@@ 4 Experiments
Table 1: Comprehensive evaluation of 2-bit weight quantization (W2A16) on LLaMA-2 models. Some results are from the LLM quantization benchmarking paper [Zhao et al., 2025].
@@ 4.1 Experimental Setup
We evaluate ButterflyQuant on LLaMA-2-7B and LLaMA-2-13B [Touvron et al., 2023], using WikiText-2 [Merity et al., 2017] and C4 [Raffel et al., 2020] for perplexity evaluation, plus six zero-shot reasoning tasks: WinoGrande [Sakaguchi et al., 2020], PIQA [Bisk et al., 2020], HellaSwag [Zellers et al., 2019], ARC [Clark et al., 2018], and MMLU [Hendrycks et al., 2021]. We compare against GPTQ [Frantar et al., 2023], AWQ [Lin et al., 2024a], OmniQuant [Shao et al., 2023], QuIP [Chee et al., 2023] and SpinQuant [Liu et al., 2024a] under aggressive 2-bit weight quantization (2-bit weights with 16-bit activations, W2A16). More experiments on other LLM architectures can be found in Appendix.
@@ 4.2 Main Results
Table 1 demonstrates ButterflyQuant’s superiority in low-bit quantization across all metrics. We achieve 2.4× lower perplexity than the best baseline (15.40 vs 36.77 for GPTQ on WikiText-2 7B, 10.24 vs 13.75 for QuIP on 13B) while AWQ fails catastrophically with perplexity exceeding $10^{5}$. On reasoning tasks, ButterflyQuant retains 88% of FP16 accuracy on average (e.g., 62.27% vs 69.06% FP16 on WinoGrande, 68.97% vs 78.07% FP16 on PIQA) while baselines retain only 65-73% of FP16 performance, with consistent improvements across both model scales validating the effectiveness of our approach.
@@ 4.3 Ablation Studies
Training Dynamics and Design Validation.
@@ A.1 Implementation Details
Optimization. Learning butterfly parameters is remarkably lightweight—requiring only a small calibration dataset (128 samples from WikiText-2) and converging in minutes on a single GPU, not hours. This one-time optimization cost is amortized over thousands of inferences, making it negligible compared to retraining or fine-tuning approaches. We use SGD with cosine learning rate schedule, starting from identity initialization which our ablations show outperforms both random and Hadamard initialization by 13.4% and 6.3%, respectively. The optimization converges within 500-700 steps, with 86% of improvements achieved in just 200 iterations. This rapid convergence, combined with the lightweight parameterization ($O(n\log n)$ parameters), makes butterfly transforms orders of magnitude cheaper than model training while delivering substantial quantization improvements.
@@
Appendix B Butterfly Transform Details
B.1 Example: 4×4 Hadamard
@@
B.2 Concrete Example: Butterfly Matrix Structure for $n=8$
@@
B.3 The Butterfly Pattern Visualization
@@
B.4 Permutation Structure and Block Decomposition
@@
B.5 Alternative Factorizations for Non-Power-of-2 Dimensions
@@
  • $80\times 64$: More uniform but requires a composite butterfly for both factors
@@ B.6 Mathematical Properties
Butterfly transforms with $O(\log n)$ layers can efficiently approximate structured orthogonal matrices and exactly represent important transforms, including Hadamard, DFT, and DCT matrices.
@@
B.7 Method Characteristics and Efficiency